#1 a)   Calculate   the   difference   in   kinetic   energy   when   a   1300   kg   car   changes   its   speed   from                       90.0   m/s   to   50.0   m/s.      

  
  
  
  
  
   b)       How   much   work   is   done   on   the   car   by   friction   force   between   the   tires   and   road?       Specify       the   sign   of   work.  

please answer both as they go together
Answer #1

a)

dKE = 0.5 m ( v^2 - u^2)

dkE = 0.5*1300* ( 90^2 - 50^2)

dKE = 3.64* 10^6 J

======

b)

workdone by frictional force = change in kE

W = - 3.64* 10^6 J

=====

Comment before rate in case any doubt, will reply for sure.. goodluck